son and fathers age totals 49, father had son at age 25, how old are they now

Answer by checkley79(3341)

LET X BE THE SON'S AGE.
LET X=25 BE THE FATHER'S AGE.
X+X+25=49
2X=49-25
2X=24
X=12 AGE OF THE SON NOW.
12+25=37 AGE OF THE FATHER NOW.
PROOF:
12+37=49
49=49